OpenStudy (anonymous):

Can someone help me solve this algebraically? (x+2)/(x-8)>(x)/(x+5)

OpenStudy (anonymous):

=(x+2)(x+5)>x(x-8) x\[\neq\]8,-5 =x^2 + 7x+10>x^2-8x =15x>-10 =x>-2/3 where x\[\neq\]8

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

i think in interval notation it is \[(-5,-\frac{2}{3})\cup (8,\infty)\]

OpenStudy (anonymous):

you cannot multiply both sides by the variable so solve these, because you don't know if it is positive or negative

OpenStudy (anonymous):

you have to start with \[\frac{x+2}{x-8}-\frac{x}{x+5}>0\] and go from there

OpenStudy (anonymous):

so i am afraid your first solution is incorrect. can you continue?

OpenStudy (anonymous):

Ok. Well, your answer does look more like the examples in the book. I'm afraid I can't continue, I really don't know what i'm doing. What would I do next?

OpenStudy (anonymous):

you have to subtract on the left. \[\frac{(x+2)(x+5)-x(x-8)}{(x-8)(x+5)}>0\] then multiply out in the numerator and combine like terms

OpenStudy (anonymous):

\[\frac{x^2+7x+10-x^2+8x}{(x-8)(x+5)}>0\] \[\frac{15x+10}{(x-8)(x+5)}>0\]

OpenStudy (anonymous):

now the zeros of the factors are at \[-5,-\frac{2}{3},8\]so you have to make 4 intervals

OpenStudy (anonymous):

\[(-\infty, -5),(-5,-\frac{2}{3}),(-\frac{2}{3},8),(8,\infty)\]

OpenStudy (anonymous):

then check on one interval by picking a number for x and substituting in the expression to see if it is positive or negative. it will change signs over each interval. if you do that you will see that the answer i wrote above is correct.

OpenStudy (anonymous):

Wow. That's wonderful! Thank you! I will check to make sure it is.

OpenStudy (anonymous):

easiest number to check is at x = 0 on the interval \[(-\frac{2}{3},8)\] if you replace x by 0 you get \[\frac{10}{-8\times 5}\] which is surely negative. you want positive because you have greater than zero, so pick the appropriate intervals

OpenStudy (anonymous):

Thank you. Did you happen to see my other question? If the total weekly cost for a company to produce its product as a function of the number of units produced is given by the equation f(x)=2x+500, write an equation to represent the average cost per unit.

OpenStudy (anonymous):

I'm suppose to write an equation and the answer someone else gave me was f(x)=2+(500/x). Does that look right?